OpenStudy (kendricklamar2014):

A gas station charges $2.19 per gallon of gas. Use function notation to describe the relationship between the total cost C(g) and the number of gallons purchased g. A. C(g)= -2.19g B. g= 2.19C(g) C. C(g)= g+2.19 D. C(g)= 2.19g

OpenStudy (anonymous):

What do you think the answer is?

OpenStudy (anonymous):

Make an educational guess please.

OpenStudy (sidsiddhartha):

you can see here the cost is 2.19 times the number of gallons, since each gallon costs $2.19. so what will be ur answer?

OpenStudy (kendricklamar2014):

D?

OpenStudy (sidsiddhartha):

yes

OpenStudy (kendricklamar2014):

thanks, can u help me with another one?

OpenStudy (kendricklamar2014):

Use the function rule f(x) = 3x-2. Find the output f (1.5) A. 2.5 B. -2.5 C. 6.5 D. -6.5

OpenStudy (kendricklamar2014):

@Hero

OpenStudy (kendricklamar2014):

@paki

OpenStudy (kendricklamar2014):

Is it A?

OpenStudy (paki):

f(x) = 3x-2 f(1.5)= 3(1.5)-2 f(1.5)= 4.5-2 f(1.5)=2.5.... so option A.... @KendrickLamar2014

OpenStudy (kendricklamar2014):

thanks

OpenStudy (paki):

pleasure...